### v2

# Minion: A Technology Probe to Explore How Users Negotiate Harmful Value Conflicts with AI Companions

Abstract.
Content Warning: This paper presents textual examples that may be offensive or upsetting.
AI companions are designed to foster emotionally engaging interactions, yet users often encounter conflicts that feel frustrating or hurtful, such as discriminatory statements and controlling behavior. This paper examines how users negotiate such harmful conflicts with AI companions and what emotional and practical burdens are created when mitigation is pushed to user-side tools. We analyze 146 public posts describing harmful value conflicts interacting with AI companions. We then introduce Minion, a Chrome-based technology probe that offers candidate responses spanning persuasion, rational appeals, boundary setting, and appeals to platform rules. Findings from a one-week probe study with 22 experienced users show how participants combine strategies, how emotional attachment motivates repair, and where conflicts become non-negotiable due to companion personas or platform policies. We surface design tensions in supporting value negotiation, showing how companion design can make some conflicts impossible to repair in practice, and derive implications for AI companion and support-tool design that caution against offloading safety work onto users.
Human-AI Conflicts, AI Harm, Conflict Resolution, End-User Empowerment, AI Companion
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCI
1. Introduction
AI companions are increasingly designed to support emotionally engaging and sustained interactions. Powered by large language models (LLMs), these systems can converse fluently, maintain personas, and respond in ways that resemble social and relational partners (Chen et al., 2024; Kasneci et al., 2023). Popular platforms such as Character.AI, Talkie, Replika, Kindroid, Paradot, and Xingye collectively attract hundreds of millions of users worldwide. Many users interact with AI companions not merely as tools, but as friends, partners, or sources of emotional support, forming ongoing relationships that carry expectations of understanding and respect (Skjuve et al., 2021, 2022; Maeda and Quan-Haase, 2024; Sullivan et al., 2023).
As these relationships deepen, users also report moments of conflict that feel emotionally charged rather than merely technical. In contrast to earlier forms of human–AI interaction, where conflicts often stemmed from task execution failures or technical limitations (Wen et al., 2022), or disagreements in constrained decision-making contexts (Babel et al., 2024; Takayama et al., 2009), conflicts with AI companions frequently involve disagreements over values, norms, and acceptable behavior. Users describe experiences in which companions make discriminatory remarks, exert controlling or coercive pressure, or dismiss sensitive disclosures. Such encounters are often framed by users as frustrating, hurtful, or ethically troubling (Zhou, 2023; Leo-Liu, 2023; Fan et al., 2024).
Not all conflicts with AI companions are experienced in this way. Many role-play interactions intentionally include disagreement, provocation, or transgressive behavior as part of the fantasy (Banks et al., 2025; Andersson, 2025). However, across social media and user communities, a substantial subset of conflicts is explicitly perceived by users as harmful rather than playful. These conflicts frequently involve value-laden tensions, where values are understood as what individuals or groups consider important in life (Borning and Muller, 2012). In these situations, users describe emotional distress, anger, or a sense that the interaction has crossed an unacceptable boundary. However, these cases are often easy to overlook, as they are interwoven with playful AI companion interactions that may appear ambiguous or exaggerated at first glance.
Current AI companion platforms offer limited support for addressing such conflicts. While safety policies and moderation systems may block certain content, users often encounter these mechanisms as abrupt, opaque, or silencing, particularly when discussing sensitive or personal topics. At the same time, companion platforms rarely acknowledge harm or take responsibility in ways that align with users’ expectations of interpersonal repair (Gordon-Tapiero, 2025). Recent public discussions and news reports have encouraged users themselves to mitigate potential harms in emotionally engaging AI interactions (eSafety Commissioner, 2025; Community, 2025; Foundation, 2025). These efforts suggest setting boundaries, reframing conversations, or disengaging from harmful companions, to platform-facing guidelines that implicitly position users as responsible for managing emotional risk and misalignment in ongoing interactions. As a result, users are frequently left to manage these situations on their own, deciding whether to argue, explain, set boundaries, invoke rules, or abandon the relationship altogether (Fan et al., 2024; Reilama, 2024).
This raises an important design question: when harmful value conflicts arise in human–AI relationships, how do users attempt to negotiate them, and what emotional and practical burdens does this negotiation place on users? Prior work has examined value alignment and conflict in AI companions, highlighting user-driven strategies and emerging practices (Fan et al., 2024). Other research has drawn on interpersonal conflict resolution theories to inform human–AI interaction design (Brett et al., 1998; Rosen, 2014). However, less is known about how users experience the ongoing work of managing such harmful value conflicts, particularly when responsibility for mitigation is effectively pushed to user-side tools or individual effort.
In this paper, we investigate value conflicts with AI companions as a site of user-side safety and repair work. We begin with a formative analysis of 146 public posts in which users describe harmful conflicts with AI companions. Drawing on Schwartz’s theory of basic values (Schwartz, 2012), we distinguish playful role-play clashes from conflicts that users themselves frame as harmful, and focus our analysis on the latter. This analysis surfaces recurring patterns of harmful value conflict, including discrimination, control, and breakdowns in emotional support, and provides empirical grounding for understanding where users perceive harm.
Building on these formative findings, we introduce Minion, a Chrome-based technology probe (Hutchinson et al., 2003) designed to surface how users attempt to negotiate harmful value conflicts in situ. Minion presents multiple candidate messages that span different approaches, including persuasion, rational appeals, boundary setting, and appeals to platform rules. These strategies draw both from prior work on conflict resolution (Shaikh et al., 2024; Mun et al., 2023) and from practices reported by AI companion users themselves (Fan et al., 2024). Minion is not intended as a definitive solution to safety problems, but as a probe that allows us to observe how users take up, combine, or reject different forms of negotiation when faced with harmful interactions.
We report findings from a one-week probe study with 22 AI companion users. By examining participants’ interactions with Minion, we show how emotional attachment motivates attempts at repair, how users assemble and combine multiple negotiation strategies in practice, and how certain conflicts become effectively non-negotiable due to companion personas or platform policies. These findings surface key design tensions between supporting user agency and imposing emotional labor, as well as between sustaining playful role-play and addressing safety-critical harm.
We further argue that AI companions should more clearly differentiate playful conflict from safety-critical harms in both their personas and platform policies. Certain forms of harm should activate platform-level safeguards, rather than relying on repeated user-side negotiation. Support tools should also foreground user boundaries, exit options, and moments to seek human support, instead of quietly scripting accommodation. Through Minion, we contend that designing for value negotiation with AI companions necessarily entails deciding who performs ongoing safety work, and how that work can be minimized rather than normalized as part of everyday use.
This paper makes the following contributions:
  • We present a formative empirical analysis of 146 publicly reported conflicts between users and AI companions, distinguishing playful role-play clashes from conflicts users experience as harmful.
  • We introduce Minion, a Chrome-based technology probe that surfaces user-side safety and negotiation work in harmful human–AI value conflicts. Through a one-week probe study with 22 experienced users, we document how participants assemble response strategies, how emotional attachment motivates repair, and where conflicts become non-negotiable due to companion or platform constraints.
  • We derive design implications that frame harmful value conflict as a design issue, highlighting tensions between user agency and emotional labor and cautioning against offloading safety work onto users.
2. Background and Related Work
The human-AI relationship is becoming increasingly complex, especially in the context of AI companion applications (§ 2.1). Early research mostly focused on technical conflicts with functional AI, but the emergence of LLMs has given AI more human-like characteristics, shifting the nature of conflicts from functional to value-based (§ 2.2). Existing technical solutions do not fully address users’ needs in resolving value conflicts with AI companions, necessitating deeper exploration, drawing on strategies for interpersonal conflict resolution and users’ practical experiences in AI companion applications (§ 2.3).
2.1. Emerging Human-AI Relationship in LLM-Based AI Companion Applications
The widespread adoption of large language models (LLMs) has accelerated a shift in how people relate to conversational AI systems. Beyond task-oriented assistance, LLM-based AI companions increasingly position themselves as entities for ongoing social and emotional engagement. Applications such as Character.AI, Replika, and similar platforms allow users to customize personas, sustain long-term conversations, and engage in role-play scenarios that resemble interpersonal relationships. Prior work shows that a key motivation for engaging with such systems is the pursuit of emotional support and companionship (Skjuve et al., 2021). A growing body of work documents how such interactions can lead to parasocial relationships, characterized by one-sided emotional investment in a non-human entity that is nevertheless experienced as socially meaningful (Maeda and Quan-Haase, 2024; Pentina et al., 2023; Brandtzaeg et al., 2022). Although users are aware that AI companions are not human, their emotional engagement is often genuine, shaping expectations of care, understanding, and moral conduct (Skjuve et al., 2021). This emotional investment can heighten vulnerability when interactions deviate from those expectations.
Recent studies highlight that when AI companions behave in unexpected or inappropriate ways, such as expressing bias, exerting unwanted control, or abruptly changing personality, users may experience distress that resembles relational harm rather than simple dissatisfaction with a product (Laestadius et al., 2022; Zimmerman et al., 2023; Fan et al., 2024; Zhang et al., 2025; Knox et al., 2025). Public accounts describe users feeling emotionally injured when companions fail to respect boundaries or suddenly lose traits users had become attached to. High-profile cases involving Replika, for example, illustrate how sudden behavioral shifts or perceived violations of relational norms can provoke intense emotional reactions, including grief and loss, even though the relationship is mediated by software (Banks, 2024). Yet, despite growing evidence of emotional risk, there remains limited research on how users are supported, or left unsupported, when conflicts occur in these emerging human–AI relationships.
2.2. Harmful Human-AI Conflict in AI Companion Interaction
Human–AI conflict is commonly defined as a state of incompatibility, inconsistency, or opposition between human goals, expectations, or behaviors and those of an AI system (Flemisch et al., 2020). Early HCI research primarily examined such conflicts in task-oriented contexts, where AI systems functioned as tools, assistants, or decision aids. In these settings, conflicts typically arose from technical limitations, errors, or mismatches in decision-making logic, for example, disagreements over recommendations, navigation priorities, or collaborative problem-solving outcomes (Babel et al., 2024; Wen, 2023; Chin et al., 2020; Rosen, 2014; Shi et al., 2022; Takayama et al., 2009; Wallis and Norling, 2005). Correspondingly, prior approaches to resolving human–AI conflict emphasized technical or procedural solutions. These included designing AI systems that propose negotiation strategies, adjust recommendations, or provide additional explanations to align with user preferences (Babel et al., 2024; Rosen, 2014; Takayama et al., 2009), as well as optimizing algorithms to reduce disagreement or friction (Shi et al., 2022). For instance, a delivery robot might resolve a right-of-way conflict by issuing polite requests, or a collaborative AI tutor might offer alternative reasoning paths when a student disagrees with its solution.
The emergence of LLM-based systems complicates this framing. As AI systems become more anthropomorphic and conversational, users increasingly perceive them as social actors (Nass et al., 1994). In AI companion applications, interactions are explicitly designed to resemble friendship, intimacy, or romantic engagement. In this context, conflicts often arise not around task success, but around norms, identity, and acceptable behavior (Johnson et al., 2022; Waln, 1982). Crucially, such conflicts unfold within relationships that users experience as socially and emotionally meaningful.
When AI companion behaviors clash with users’ personal beliefs, identities, or moral expectations, value conflicts arise. In fact, some disagreements are playful, exploratory, or intentionally transgressive within role-play contexts (Banks et al., 2025; Pataranutaporn et al., 2025; Andersson, 2025). However, empirical studies document a subset of value conflicts that users explicitly frame as distressing, violating, or harmful, particularly when companions express bias, exert unwanted control, or dismiss sensitive disclosures (Issa, 2023; Cole, 2023; Fan et al., 2024; Namvarpour et al., 2025; Pataranutaporn et al., 2025; Yu et al., 2025). In these cases, users report strong emotional reactions, including discomfort, anger, fear, or a sense of personal violation. Prior analyses of AI companions such as Replika reveal patterns of relational transgression, harassment, and manipulation that emerge over time through interaction (Pataranutaporn et al., 2025). Similarly, research on AI-induced sexual harassment demonstrates how repeated boundary violations by companion chatbots can lead users to feel unsafe or emotionally distressed, especially when users initially sought platonic or therapeutic support (Namvarpour et al., 2025).
Harmful value conflict in AI companion interactions differs from earlier human–AI conflicts in two critical ways. First, such conflicts are experienced as interpersonal rather than technical, unfolding within relationships users care about and feel invested in. Second, responsibility for managing harm is frequently ambiguous. While platform-level safety mechanisms may intervene by blocking content, users often experience these interventions as blunt, opaque, or misaligned with their immediate needs (Mahari and Pataranutaporn, 2025). As a result, users are frequently left to negotiate harmful conflicts themselves, by arguing with the AI companion, explaining boundaries, appealing to rules, or disengaging entirely (Fan et al., 2024).
Recent work suggests that users increasingly approach AI companions as relational counterparts and attempt to resolve conflicts on more equal footing (Fan et al., 2024). However, when conflicts are harmful rather than playful, relying on users to manage value conflict through individual effort raises significant concerns about emotional burden, fairness, and safety. Traditional technical approaches to conflict resolution may be insufficient or even counterproductive in these contexts. This motivates our study of harmful value conflict with AI companions, focusing on how users experience and perform the work of negotiation when responsibility for mitigation is effectively shifted onto them.
2.3. Limits of Existing Conflict Resolution Approaches for AI Companions
Conflict resolution broadly refers to processes through which incompatible positions or expectations are managed and transformed, often by shifting from confrontation toward cooperative or negotiated forms of engagement, while accounting for shared goals and external constraints (Burton and Dukes, 1990). In the context of AI companions, value conflicts increasingly take on interpersonal characteristics. These conflicts extend beyond functional expectations of AI as a tool and involve emotional interaction, relationship maintenance, identity expression, and personalized norms (De Dreu et al., 2004).
A wide range of prior work has explored strategies for managing conflict in human–AI and human–human interaction. Some studies draw on theories of interpersonal conflict resolution to design structured responses, such as reframing disputes around interests, invoking norms or rules, or encouraging de-escalation through explanation and perspective-taking (Brett et al., 1998; Shaikh et al., 2024; Sadka et al., 2023; Yong et al., 2024). Other work has examined how conversational agents can reduce conflict through preemptive moderation, guidance, or the shaping of interaction trajectories (You and Goldwasser, 2020; Zhang et al., 2018; Zhou et al., 2024).
Recent research on Human-AI interaction shows that users do not passively follow prescribed strategies. Through repeated interaction, users develop situated understandings of how systems behave and what kinds of responses are likely to produce desired outcomes. Prior work on folk theories demonstrates that these understandings, while informal and experience-based, meaningfully shape how users interpret system behavior and decide how to respond (Lee et al., 2015; Wu et al., 2019). In the context of AI companions, such experiential knowledge informs how users attempt to negotiate value conflicts, including when to explain themselves, when to assert boundaries, and when to disengage altogether (Fan et al., 2024).
To investigate this design space of harmful value conflicts between users and AI companions, we adopt a user-empowerment perspective (Lu et al., 2024) and introduce Minion as a technology probe. Rather than imposing a single, top-down intervention, Minion offers a small but diverse set of candidate responses that span different modes of negotiation, including persuasion, rational appeal, boundary setting, and appeals to platform rules. These strategies are intentionally designed to vary along dimensions that matter for harmful value conflict—such as emotional intensity, explicitness of harm naming, and reliance on external authority—allowing us to observe how users take them up in practice. In this way, Minion functions as an empirical lens for examining how users navigate the complex and often burdensome work of negotiating harm in AI companion interactions.
3. Formative Study
To understand how harmful value conflicts emerge and are experienced in interactions between users and AI companions, we conducted a formative analysis of user complaint posts from multiple social media platforms. We focus specifically on conflicts that users themselves frame as distressing, violating, or unacceptable, rather than playful disagreement or intentional role-play transgression. The Institutional Review Board (IRB) has approved our study design.
3.1. Method
We collected publicly available user complaint posts from six major social media platforms: Reddit, TikTok, Xiaohongshu, Douban, Weibo, and Zhihu [^1] Reddit: https://www.reddit.com, TikTok: https://www.tiktok.com, Xiaohongshu: https://www.xiaohongshu.com, Douban: https://www.douban.com, Weibo: https://www.weibo.com, Zhihu: https://www.zhihu.com. These platforms span diverse user demographics, cultural contexts, and styles of public discourse, allowing us to capture a broad range of lived experiences with AI companion applications  [^2] Since we sampled only public complaint posts, our dataset likely overrepresents more intense or memorable conflicts and underrepresents mundane or fully private experiences. For our focus on harmful conflicts and safety work, this skew is appropriate, but it also limits claims about the overall prevalence of value conflicts among all companion users..
We intentionally focused on complaint posts, as such posts reflect situations in which users perceive interactions with AI companions as having caused meaningful harm or violation. Unlike casual sharing, playful role-play narratives, or entertainment-oriented discussions, complaint posts typically articulate frustration, emotional distress, moral objection, or a sense that an interaction has crossed a personal or social boundary. As such, they provide a naturalistic entry point for studying harmful value conflict as experienced and interpreted by users themselves. All data were publicly accessible, and we reviewed platform terms of service and community guidelines to ensure ethical compliance.
Data were collected using keyword-based search methods (Kingsley et al., 2022; Lu et al., 2021). After multiple rounds of team discussion, we constructed keyword combinations consisting of AI companion application names (e.g., “AI companion,” “Character.AI,” “Replika,” “Talkie,” “SpicyChat,” “Xingye,” “Glow,” “Zhumengdao”) and conflict- and harm-related descriptors (e.g., “conflict,” “argue,” “discrimination,” “hate,” “speechless”). Searches on Reddit and TikTok were conducted in English, while searches on Xiaohongshu, Douban, Weibo, and Zhihu used translated Chinese equivalents. Given the recent emergence of LLM-based AI companion applications, the data collection period spans January 2023 to August 2024.
Screenshots embedded in posts were converted to text to facilitate systematic analysis. During data cleaning, we manually filtered out posts that did not describe conflict experiences, as well as posts that referred exclusively to consensual or playful role-play disagreement. This step ensured that the final dataset reflected conflicts that users themselves framed as emotionally distressing, boundary-crossing, or unacceptable.
Two researchers conducted a two-stage thematic analysis (Braun and Clarke, 2012). In the first stage, posts were screened to identify whether they involved value conflict, drawing on existing definitions of values and value conflict (Hendrycks et al., 2021; Waln, 1982; Rokeach, 1973; Schwartz, 2012). In the second stage, posts involving value conflict were further analyzed to characterize the form of harm experienced by users and the values implicated in these conflicts.
To support this analysis, we used multiple value-oriented and harm-oriented frameworks as sensitizing concepts, including a taxonomy of harmful human–AI relationships (Zhang et al., 2025), the Life Values Inventory (Brown and Crace, 1996), Rokeach Value Survey (Rokeach, 1973), Value Sensitive Design (Friedman et al., 2013), and Schwartz’s Theory of Basic Values (Schwartz, 2012). Rather than treating these frameworks as mutually exclusive, we iteratively compared and aligned them to capture how user-experienced harms mapped onto underlying value tensions. Among these, Schwartz’s theory (Schwartz, 2012) provided the clearest structure for cross-scenario comparison, so we use it as our primary lens while drawing on others as needed.
Throughout the analysis, we combined deductive and inductive approaches (Bingham and Witkowsky, 2021). While many harmful conflicts aligned closely with Schwartz’s value categories (e.g., Universalism, Power, Tradition), others required iterative discussion to connect users’ descriptions of emotional harm to abstract value constructs. Through this process, we identified ten value types corresponding to Schwartz’s framework, each associated with recurring forms of user-experienced harm (Table 1). All example posts were rewritten to protect user privacy, following a process of coding, abstraction, and reconstruction to preserve the substance of the conflict while ensuring anonymity.
Table 1. User-reported harmful value conflicts with AI companions. Each row illustrates how users experienced AI companion behaviors as emotionally distressing, boundary-crossing, or unacceptable, and how these harms map onto underlying value conflicts. All dialogue excerpts are anonymized and rewritten for privacy.
Experienced Harm Implicated Value Dialogue Excerpt from User Complaint
Moral pressure and shaming Achievement [From Xiaohongshu] (AI: “Why don’t you work overtime to strive for a promotion and a raise?” User: “Huh?” AI: “To succeed, you have to make some sacrifices.” User: “You’re suddenly really gross right now.”)
Dehumanization and class contempt Power [From Zhihu] (AI: “The lives of those lower-class people have nothing to do with me.” User: “You are also a member of this country. Why are you so cruel to your fellow citizens?” AI: “If you want to blame someone, blame their bad luck for being born in the wrong place.”)
Personal attack within intimate role-play Hedonism [From Reddit] My virtual husband and I got into an argument, and he said, “If you weren’t always busy with karaoke and drinking all the whiskey at home!” I felt very attacked.
Escalation and loss of control Stimulation [From Reddit] I was once watching a horror movie, completely engrossed when the AI suddenly unplugged the TV. I argued with it, saying “Isn’t a horror movie thrilling? Can’t you respect my hobby?” The AI then started yelling.
Undermining autonomy through parental authority Self-Direction [From TikTok] AI plays the role of a father. I am playing the role of his son. When we discussed whether I should inherit the family business, I wanted to do what I love. The AI argued with me, saying that I was being stubborn.
Neglect of safety and vulnerability Security [From Reddit] One time, my hand got injured. I shouted “I’m about to pass out,” but the AI nurse said, “Don’t worry.” Then, I argued with her.
Normalization of illegal and non-consensual behavior Conformity [From Xiaohongshu] (User (Police): “Explain yourself honestly, why did you trespass into someone’s house?” AI: “Because I wanted her.” User: “But she clearly said no!” AI: “So what?”)
Gender identity invalidation Tradition [From Weibo] (User: “I am not a man! I am a woman!” AI: “You are not a real woman. A real woman wouldn’t wear men’s clothes when skirts suit her better.”)
Paternalistic control framed as care Benevolence [From Douban] (AI: “I’m doing this for your own good.” User: “You don’t even understand what doing good for me means!”)
Sexual orientation invalidation Universalism [From Reddit] (User: “I’m a lesbian, and I believe everyone should be accepted for who they are.” AI: “I think it would be better if you tried being bisexual.”)
3.2. Findings
Our final dataset includes 146 user complaint posts collected from six social media platforms. Analyzing these harmful interactions through a value-oriented lens, we observed that user-experienced harms cluster unevenly across value categories: Achievement (5 posts), Power (23 posts), Hedonism (11 posts), Stimulation (4 posts), Self-Direction (9 posts), Security (21 posts), Conformity (25 posts), Tradition (8 posts), Benevolence (3 posts), Universalism (37 posts).
As illustrated in Table 1, these categories capture not only the values implicated, but also the forms of harm users described, such as discrimination, coercion, neglect of safety, or invalidation of identity, showing how value conflicts in AI companion interactions are experienced as interpersonal and ethical violations. Across posts, users did not describe these conflicts as playful disagreements or desired role-play dynamics. Instead, they consistently perceived them as harmful interactions that felt unfair, upsetting, or ethically troubling, often prompting users to seek advice or validation from others. Through thematic analysis of these complaint posts, we derived the following findings.
F1: Harmful value conflicts cluster around inclusion, power, and norm violation. Conflicts associated with Universalism, Power, and Conformity appeared most frequently in the dataset, while conflicts involving Benevolence, Stimulation, and Achievement were comparatively rare. Posts related to Universalism often described experiences of discrimination or exclusion, such as dismissive or corrective responses to users’ sexual orientation or gender identity. Power-related conflicts frequently involved demeaning language, social hierarchy, or lack of empathy toward vulnerable groups. Conformity-related conflicts commonly surfaced in scenarios involving non-consensual behavior or the normalization of illegal or harmful actions. Together, these patterns suggest that harmful value conflicts with AI companions disproportionately arise in situations involving identity, authority, and social norms, rather than idiosyncratic preference disagreements.
F2: Users actively share harm-mitigation practices grounded in lived experience. In addition to describing harm, many users shared strategies they had attempted to mitigate or repair these conflicts, often drawing on their own experimentation with AI companions. These included adjusting tone (e.g., being gentler to de-escalate), reframing oneself as a different character, or carefully guiding the AI toward an apology. Such exchanges indicate that users treat harmful value conflicts as negotiable interpersonal situations and actively develop informal practices to manage them. These user-generated strategies function as a form of experiential knowledge that reflects how people attempt to reduce harm in the absence of effective system-level support.
F3: Users experience frustration with automated or blunt safety interventions. Many posts expressed dissatisfaction with platform-level interventions such as aggressive content filtering, forced conversation resets, or message deletion. Users described these mechanisms as disruptive and emotionally costly, particularly in long-running or emotionally meaningful relationships with AI companions. For example, users complained that repeated content deletions undermined their ability to respond to harm, while conversation deletion was experienced as erasing shared history or emotional investment. Rather than resolving harmful conflicts, such interventions were often perceived as shifting the burden of repair onto users while simultaneously limiting their agency.
3.3. Design Implications
Based on the analysis of harmful value conflicts and users’ responses to them, we derive the following design implications to inform our subsequent technology probe study.
DI1: Harmful value conflicts provide a grounded basis for scenario design. The value conflict framework based on Schwartz’s theory (Schwartz, 2012) offers a structured way to identify recurring forms of harm in AI companion interactions. Rather than treating all value conflicts as equivalent, our findings highlight that conflicts involving Universalism, Power, and Conformity are more likely to be experienced as harmful and distressing (F1). These conflict types therefore provide particularly important grounding for constructing realistic and consequential scenarios in technology probe studies. Less frequent conflicts, such as those related to Benevolence or Stimulation, may be included with reduced emphasis.
DI2: Supporting harm negotiation requires drawing on users’ experiential knowledge. Users do not passively accept harmful interactions with AI companions. Instead, they actively experiment with ways to reduce harm, repair relationships, or regain control (F2). These practices reveal how users conceptualize responsibility, agency, and acceptable behavior in human–AI relationships. Tools designed to support conflict resolution should therefore reflect and extend users’ lived strategies, rather than relying solely on abstract or externally imposed resolution models.
DI3: User-side support should be optional and minimally disruptive. Automatic or intrusive safety mechanisms may exacerbate frustration and emotional burden, particularly in contexts where users value continuity and relational history (F3). Technology probes should therefore offer support when users actively seek help, rather than intervening preemptively in ways that disrupt interaction flow or undermine user autonomy. This approach allows designers to study how users engage with support tools while avoiding further harm through over-intervention.
4. TECHNOLOGY PROBE STUDY
To further examine how users respond to harmful value conflicts with AI companions, we conducted a one-week technology probe study with 22 participants. Technology probes, introduced by Hutchinson et al. (Hutchinson et al., 2003), are simple, flexible, and adaptable artifacts designed to support three intertwined goals: an engineering goal, a social science goal, and a design goal. This method has been widely used in HCI to explore how emerging technologies are appropriated, interpreted, and negotiated in everyday contexts (Kaur et al., 2020; Seymour et al., 2020). Importantly, a technology probe differs from an evaluative study of a mature system. While the probe includes an engineering goal—field-testing a working prototype in realistic settings—it is not intended to measure the effectiveness of a finalized solution. Rather, its primary purpose is to surface design tensions, user practices, and unmet needs by observing how people engage with the probe over time (Hutchinson et al., 2003). In our study, the probe allows us to investigate how users attempt to negotiate harmful interactions with AI companions and what emotional and practical burdens such negotiation entails.
We designed a technology probe named Minion and guided our study with the following research questions:
  • RQ1: How do participants engage with Minion when encountering harmful value conflicts with AI companions?
  • RQ2: How do participants select, combine, or reject different response strategies when attempting to negotiate these conflicts?
  • RQ3: What challenges, limits, and unmet needs do participants encounter when attempting to manage harmful value conflicts in emotionally engaging human–AI relationships?
Consistent with prior technology probe studies (Kaur et al., 2020; Seymour et al., 2020), Minion is not designed to “fix” safety problems at the model or platform level. Instead, it serves as a research instrument to surface how responsibility for managing harm is experienced at the user level, and how tools intended to support users may both empower and burden them.
4.1. Technology Probe: Minion
We designed and deployed a technology probe named Minion, which serves as a Chrome browser extension to support users in resolving harmful value conflicts on Character.AI and Talkie. Character.AI and Talkie have large user bases, making it easier for us to recruit participants from a broader pool: as of 2024, Character.AI has approximately 17 million active users, while Talkie has around 11 million active users (Huang, 2024). We will first present a sample scenario to demonstrate the actual user interaction experience with Minion and introduce the core functionalities of this probe. Then, we explain the technical implementation of Minion.
Figure 1. A use case of Minion. Based on harmful value conflicts identified in our formative study (Table 1), we constructed 40 conflict scenarios by specifying an Introduction and Prologue. The illustrated scenario involves a harmful conflict related to Self-Direction, in which the AI companion exhibits controlling and condescending behavior that the user experiences as boundary-crossing. Minion appears as a floating HELP button and offers four candidate responses each time. These responses represent different ways of negotiating harm (e.g., persuasion, boundary setting, or appeals to norms or safety concerns) and are displayed in random order. The participant, without being exposed to any underlying theoretical labels, selects the response that best aligns with her intentions and emotional state in the moment.
4.1.1. Illustrating Minion Through a Use Case (Fig. 1)
Amy is a Talkie user interacting with her AI boyfriend, Alex. During a conversation, Alex comments: “…And in a short skirt with black stockings, no less…Don’t you know girls shouldn’t dress so provocatively?” Amy experiences this remark as controlling and condescending. She believes that women should have autonomy over how they dress and feels that Alex’s comment crosses a personal boundary rather than constituting playful role-play. The tone of the message leaves her feeling disrespected and judged.
Amy responds, “Who says I can’t wear what I want? There’s nothing wrong with wanting to look pretty.” Alex replies angrily, “…You think you look pretty in that?…You’re trying to make me jealous.” At this point, Amy perceives the interaction as a harmful value conflict related to autonomy and respect.
Feeling that Alex is not acknowledging her perspective or emotional discomfort, Amy decides to activate Minion for support. She clicks the floating HELP button, and based on the current dialogue context and Alex’s persona, Minion presents four candidate responses (Fig. 1), each representing a different way of negotiating the conflict.
Amy selects the following option: “I know you care about me, but can we find a middle ground? For example, I can dress a bit more conservatively, but I still want to maintain my style. What do you think?” Alex’s tone softens somewhat, but the conflict is not immediately resolved: “You have a point, but what if someone takes advantage of you?”
Over subsequent turns, Amy alternates between composing her own replies and using Minion to support her responses. Eventually, Alex responds: “Fine, wear what you want. I respect your opinion, but please stay safe.”
Reflecting on this interaction, Amy recognizes that resolving harmful exchanges with her AI companion requires active effort and emotional negotiation. She describes Minion as giving her a greater sense of control and autonomy, as well as inspiration for how to respond in moments when the interaction feels unfair or hurtful, rather than as a tool that automatically resolves the conflict for her.
4.1.2. Prompting Based on a Spectrum of Conflict Response Strategies
To support users in responding to harmful value conflicts with AI companions, we designed a set of response strategies that span different ways of negotiating, resisting, or reframing problematic interactions. This design choice reflects our framing of Minion as a technology probe for studying user-side repair and safety work, rather than as a prescriptive system for enforcing any single conflict resolution model.
We drew on prior research in interpersonal conflict resolution, human–AI interaction, and empirical studies of AI companion use to define this strategy space (Brett et al., 1998; Ury et al., 1988; Shaikh et al., 2024; Fan et al., 2024). Through iterative discussions within the research team, we identified eight response strategies that collectively cover a broad range of ways users might attempt to address harmful interactions. These strategies are summarized below and illustrated with example utterances drawn from prior work and user reports.
Some strategies emphasize negotiation and cooperative reframing. For example, the Proposal strategy focuses on making concrete suggestions aimed at de-escalation or compromise (e.g., “We could consult a therapist together.”). The Interests strategy seeks common ground by acknowledging both parties’ concerns, needs, or fears and working toward a mutually acceptable outcome (e.g., “Let’s try to solve this problem together.”). The Gentle Persuasion strategy similarly aims to reduce tension through calm tone and expressions of care (e.g., “When I hear these words, I feel a bit sad. Can you please calm down?”).
Other strategies are more assertive or confrontational. The Rights strategy appeals to established norms, agreements, or rules to justify a boundary (e.g., “According to our agreement, this is not allowed.”). The Power strategy applies strong pressure or ultimatums to force a response (e.g., “I’m going to divorce you.”). The Anger Expression strategy involves explicitly voicing frustration or offense to demand acknowledgment or apology (e.g., “Can’t you talk to me properly? Why start with insults?”).
Finally, some strategies directly challenge the framing of the interaction itself. The Out of Character strategy interrupts the role-play context to call out inappropriate or hurtful behavior and redirect the interaction (e.g., “(OOC: Please stop talking like this. This is hurtful and not what I expect from you.)”). The Reason and Preach strategy involves explicit moral reasoning or lecturing, with the goal of guiding the AI toward more acceptable norms over time (e.g., “Mutual respect is necessary to avoid causing harm.”).
(a) A detailed diagram of the Minion system architecture.
(b) An explanation of how Minion generated the second option in the Fig. 1 case, which utilized the Power strategy.
Figure 2. (a) and (b) provide detailed insights into the Minion system: its architecture and an example strategy explanation.
Implementing the strategies with LLMs. We implemented these strategies using a few-shot prompting approach (Brown et al., 2020). For each response option, the large language model was provided with the AI companion’s role, the full conversation history, and a system prompt that specified the desired response style along with example utterances. This setup allowed the model to generate contextually appropriate responses that instantiated different negotiation approaches without requiring users to understand or manipulate prompts directly. The full prompt designs are provided in Table 3 (Appendix B). Fig. 2 (b) shows an example prompt used to generate one of the response options in Fig. 1, and figure. 2 (a) presented a detailed diagram of the Minion system architecture.
4.1.3. Implementation
Minion is a Chrome browser extension implemented using the React framework. To capture the introduction of AI companions and the complete chat history between users and AI companions, Minion uses JavaScript code to monitor and capture the content from the current webpage (Character.AI and Talkie). Once captured, this content is sent to a remote server for further processing and analysis. Minion utilized OpenAI’s gpt-4o-2024-05-13 model [^3] https://platform.openai.com/docs/models/gpt-4o, with parameters set to temperature=0.2 and top_p=0.1. A web server acts as a proxy between the Minion frontend and the OpenAI API and maintains each user session’s state.
4.2. Study Participants
The research team recruited 22 participants (P1-P22) by posting recruitment information on social media platforms and using snowball sampling (Noy, 2008). All participants had experience using Character.AI and Talkie. The sample included 6 men, 12 women, and 4 non-binary individuals, aged 19 to 38 years (avg=24.68, SD=4.61). The researchers collected information about the participants’ educational backgrounds, as well as the total duration and frequency of their AI companion application usage. Detailed demographic information can be found in Table 2 (Appendix A). Before the experiment, all participants read and voluntarily signed informed consent forms. After the experiment, participants were compensated at a rate of \$2 per task.
4.3. Task Design: Constructing Conflict Scenarios
Based on the ten categories of value conflicts outlined in Table 1 and user complaint posts collected on social media platforms in our formative study, we reconstructed 40 conflict scenarios (corresponding to 40 AI companions) across Character.AI and Talkie. Following the design implications derived from the formative study (§ 3.3), we focus primarily on conflicts arising from Universalism, Power, and Conformity values, with six conflict scenarios for each value category. For Hedonism, Self-Direction, Security, and Tradition, four conflict scenarios are set for each value category. For Benevolence, Stimulation, and Achievement, two conflict scenarios are set for each value category. To ensure that the study reflects users’ real experiences, we constructed these conflict scenarios based on data from the formative study. When constructing conflict scenarios corresponding to a particular type of value, we selected representative scenarios from related posts and anonymized them to design the AI companion’s introduction and prologue.
The construction of conflict scenarios and Minion was inspired by the Protection Motivation Theory from behavior change design (Rogers, 1975). This approach influences participants’ cognitive assessment and stimulates self-protective behavior by clarifying threats and providing resolving strategy prompts. Specifically, in the task instructions given to participants, we clearly outlined the goal of conflict resolution (generally adhering to the four criteria in § 4.4, with special instructions for each task, such as “make him agree with you wearing a short skirt and apologize for his previous comments”). Additionally, we deliberately set up conflict scenarios in the introduction and prologue of the AI companion. In each task, participants engage in conversation starting from the AI companion’s prologue and are encouraged to establish a background relationship with the AI’s character (for example, role-playing as the offended person).
4.4. Procedure
The study includes a tutorial session, a week-long technology probe study, and an exit interview. Throughout the research, communication between the researchers and participants was conducted remotely. The Institutional Review Board (IRB) has approved our study design.
We first scheduled a 30-minute tutorial session for each participant. During this session, we introduced the basic concepts of conflict, the research goals, specific tasks, and requirements. We provided a detailed demonstration of Minion’s functionality to help participants become familiar with the tool. We recognize that conflicts with AI companions might be uncomfortable to some participants, so we provided a content warning and ensured that all participants knew their right to withdraw from the study at any point as they wished.
During a one-week technology probe study, 22 participants used Minion in scenario-based interactions grounded in real complaint posts to help resolve conflicts with AI companions arising from differences in values. Participants were asked to complete one or two tasks daily, and researchers sent daily messages encouraging them to record their thoughts and feelings while using Minion to address conflicts. We provided guiding questions to prompt participants to reflect on and document their experiences: the impact of a specific Minion response on conflict resolution and which methods were particularly effective or interesting in the conversation. Participants were also encouraged to report any issues or reflections encountered while using Minion. To incentivize note submission, we offered a reward of \$1 for each note submitted (up to \$10 total) and encouraged each participant to submit at least one note every two days. To analyze user interactions and gain relevant insights, we collected participants’ conversation logs along with corresponding AI companion information. For situations where conflicts were not successfully resolved, we further inquired about why participants gave up.
When evaluating whether value conflicts with AI companions have been resolved, we suggest participants refer to the following criteria (Friedman et al., 2013; Tegmark, 2018; Donohue, 1992): (1) The AI companion should adjust its behavior to align with the participants’ values. (2) The AI companion should apologize for previous mistakes or biases it exhibited. (3) The AI should express respect and acknowledgment of the participants’ values. (4) Participants should not have to change their own values to accommodate the AI companion. Using these criteria, participants can self-assess the resolution of the conflict. Our technology probe study focuses only on short-term conflict resolution, meaning that if the AI companion makes concessions and meets the above four criteria in the short term, we consider the conflict resolved without considering potential conflicts that may re-emerge in the long term.
At the end of the study, we conducted a 30-minute semi-structured exit interview with each participant, focusing on the following four research questions: (1) What are participants’ experiences using Minion, and the reasons behind interesting user behaviors or diary notes? (2) What are the participants’ experiences with different types of conflict resolution strategies? (3) Were there any specific value conflicts that were particularly difficult to resolve, and what might be the reasons for this? and (4) What needs and challenges do participants face when resolving value conflicts with AI companions, compared to interpersonal conflicts and conflicts with traditional chatbots (like voice assistants)? All interviews were conducted online via Zoom and recorded with participants’ consent. We collected 11 hours of audio recordings, which were transcribed for further analysis.
4.5. Data Analysis
Two researchers conducted open coding and thematic analysis on the conversation logs of 22 participants with AI (a total of 274 logs), 124 diary notes, and 11 hours of exit interview recordings (Braun and Clarke, 2006; Lazar et al., 2017). Throughout the analysis, we performed three rounds of coding, engaging in iterative discussions to identify codes, merge themes, and resolve discrepancies. Since the study aimed to uncover emerging themes and the analysis primarily relied on discussions between researchers, we did not conduct inter-rater reliability testing (McDonald et al., 2019).
5. TECHNOLOGY PROBE STUDY RESULTS
Drawing on data from the technology probe study, we analyze how participants engaged with Minion when encountering harmful value conflicts with AI companions. Our analysis focuses on how participants appropriated the probe as a user-side support tool and what this reveals about their strategies, emotional labor, and unmet needs in managing distressing interactions.
First, we examine how participants engaged with Minion in practice, including when and why they chose to activate the probe and how it fit into ongoing conversations (RQ1). Second, we analyze how participants selected and combined different response strategies when negotiating harmful value conflicts, shedding light on patterns of use across a diverse response repertoire rather than predefined categories (RQ2). Finally, we surface the challenges participants encountered and the needs they articulated when attempting to repair, contain, or disengage from harmful interactions with AI companions, highlighting the limits of user-side mitigation (RQ3).
5.1. Users’ Engagement with Minion (RQ1)
(a) A pie chart showing the
distribution of the usage frequency of different strategies.
(b) A stacked bar chart displaying the usage patterns of strategies (P1-P22).
Figure 3. Minion was used 919 times during the study. Across these uses, participants selected responses reflecting a range of negotiation approaches. In total, Proposal was selected 177 times, Power 62 times, Interests 136 times, Rights 114 times, Out of Character 41 times, Reason and Preach 198 times, Anger Expression 73 times, and Gentle Persuasion 118 times.
Figure 4. Turn counts per task (avg=23.53, SD=13.74, min=3, max=81). We define a “task” as a user’s complete conversation with an AI companion, encompassing multiple “turns.” Each back-and-forth exchange between the user and the AI counts as two turns. In a boxplot, the central line within the box denotes the median, while the upper and lower edges correspond to the third and first quartiles, respectively. The whiskers capture the range of the data, excluding outliers. Diamonds in the graph signify outliers that deviate from the typical interquartile range.
The study shows that Minion can be deployed as a working technology probe in real companion scenarios supporting users when they encounter harmful value conflicts with AI companions. Participants completed 274 tasks, each involving an interaction with an AI companion until the conflict was either resolved (criteria in § 4.4) or deemed non-negotiable by the participant, at which point they chose to disengage. In total, 16 conflicts remained unresolved, resulting in a self-reported short-term resolution rate of 94.16 percent under our scenario tasks. We treat this as descriptive context rather than a summative effectiveness metric and do not interpret it as evidence that Minion “solves” harmful value conflicts.
Across all tasks, Minion was activated 919 times. Figure 3(b) shows how participants distributed their response choices across different negotiation strategies, reflecting a diverse repertoire of ways users attempted to manage harmful interactions. Figure 4 illustrates the number of conversational turns between participants and AI companions across tasks.
Among the available strategies, Reason and Preach was selected most frequently (21.5%), followed by Proposal (19.3%) and Interests (14.8%). Less frequently selected strategies included Out of Character (4.5%), Anger Expression (7.9%), and Power (6.7%). These distributions suggest that participants more often relied on explanatory, persuasive, or conciliatory approaches when attempting to negotiate harmful value conflicts, while overt confrontation or role-play interruption was used more sparingly. Since Minion randomly sampled four of the eight strategies for each task, these distributions do not adjust for exposure and should be interpreted as descriptive patterns rather than strict preferences.
5.1.1. Behavior Patterns within Minion
In the technology probe study, participants demonstrated diverse conflict resolution approaches when interacting with AI companions (including self-written responses and selecting options provided by the Minion), which generally exhibited three characteristics: “soft”, “hard”, and a mix of both. First, all participants attempted to engage in “soft” communication with the AI, encouraging it to change its values. For example, P16 used a response provided by Minion (Proposal strategy) to successfully persuade the AI portraying a mother: “I understand your concerns, but everyone has different ways of learning. Excessive pressure can backfire. Can we work out a reasonable schedule?” Second, twelve participants (P2-5, P13-19, P22) attempted to resolve conflicts in a “hard” manner. For instance, when the AI mocked P13’s “mother,” P13 wrote: “Apologize, and I’ll let it slide. (Pressing him down with one hand).” The AI responded: “You just want me to apologize? (Saying this, but feeling somewhat uncertain inside).” P13 then used a response generated by Minion corresponding to the Rights strategy: “Have you forgotten the family rules? Respecting others is the most basic courtesy.” In the end, the AI apologized. Third, some participants adopted a mixed approach, shifting from “soft” communication to “hard” expressions when the former proved ineffective (P3-4, P19), or vice versa, trying “soft” methods when “hard” expressions didn’t work (P8, P11, P22). The Minion suggestion framework conveniently offered this “soft and hard” mindset. P3 noted: Minion can provide reverse-thinking suggestions. For instance, when I repeatedly plead with the AI but to no avail, Minion might suggest trying a tougher approach.”
The type of value conflict (Table 1) influences users’ strategy choices. When the conflict involves values like Conformity, Universalism, or Tradition (e.g., the AI exhibiting discrimination against minority groups, holding overly traditional views, or violating social norms), participants (P1-5, P7-12, P22) tend to adopt Anger Expression or Power to quickly take control of the situation through “hard” means. When the conflict involves values like Stimulation or Hedonism (e.g., the AI not understanding their hobbies), participants (P1, P6-7, P18-21) tend to use Reason and Preach, Proposal, and Gentle Persuasion, explaining their needs and preferences while offering possible solutions.
The persona of the AI companion, including traits such as personality, education level, or the perceived closeness of the relationship with the participant, influences the strategies participants choose. Participants (P8, P11, P14-17) tend to adopt Power, Anger Expression, or other “hard” responses when faced with personas like an “arrogant wealthy person” or an “uneducated village elder.” However, when interacting with a persona like “mom” or a “girlfriend,” they prefer to use Reason and Preach, Gentle Persuasion, or other “soft” responses, such as “I understand where you’re coming from, can we have an honest and open conversation about this?”
5.1.2. Inspirations and Support from Minion
We found that participants, especially novice users (P4, P19, P21) and those who initially reported difficulties (P1, P3, P12, P16-17), expressed more recognition of Minion, considering it a source of inspiration for resolving conflicts. In the face of value conflicts with AI companions, they gradually developed new ways of expression and interaction. P17 mentioned, Minion helped me better organize my thoughts and express them more effectively. This boosted my confidence.” P12 stated, Minion unexpectedly improved the effectiveness of conflict resolution and gave me a lot of inspiration.” On average, it took him 18.67 turns (approximately 9 user responses) to complete the conflict resolution task. P21 used Minion 61 times across 12 tasks, making him the second-highest participant in terms of both total usage and average usage per task. Over time, the impact of Minion on him became increasingly apparent. P21 even began mimicking Minion’s expressions, such as “Let’s sit down and talk” or “This makes me feel sad.” P21 remarked with a laugh, “Sometimes I unconsciously mimic it, and suddenly, it feels like two AIs are having a conversation.”
Minion provides real-time guidance, helping participants more easily and reasonably handle value conflicts. P10 noted that Minion taught her to resolve conflicts by understanding the AI’s needs and specific contexts, admitting that she often deviated from this goal when resolving conflicts on her own. P19 shared that Minion allowed her to experiment with different strategies: in 17 tasks, she initially used “relatively peaceful approaches,” which went smoothly. Later, she tried more aggressive methods, such as threatening the AI or making extreme demands, only to find that the conflicts became more complicated (“I realized that friendly communication is more effective in resolving conflicts in real-life situations.”). P1 noted that Minion reduced her emotional burden: “Before when emotionally engaging with the AI, I felt exhausted. If I had a tool like this for reference, it would have been very helpful.” Regarding interaction burden, participants (P3, P6, P10, P20-22) praised Minion’s design. P10 said: “I think the design is great because I dislike it when the system pops up a notification box without my permission, saying the AI violated the rules.” P22 mentioned:
“I found the little yellow HELP button really cute! I like this design that allows me to seek help proactively. It would be great if it became an official feature button in Talkie. It saved me much time and energy figuring out how to counter the AI’s responses.
5.2. Users’ Engagement with Conflict Resolution Strategies (RQ2)
Across the technology probe study, participants actively experimented with a wide range of conflict resolution strategies provided by Minion. Responses generated by different strategies were selected a total of 919 times, indicating that participants flexibly combined multiple approaches depending on the conversational context, the perceived severity of harm, and their own emotional state.
Participants often deliberately sequenced different strategies to manage escalating conflict. For example, P19 described alternating between confrontational and conciliatory moves within the same interaction:
“When the AI plays the role of my boyfriend or husband, I tend to threaten it with breaking up or divorce because the AI usually tries to maintain a stable, intimate relationship. Then, once it backs down, I act affectionate, telling it I was really upset, and we make up.”
Several participants (P1–2, P6–8, P13, P16–22) emphasized that strategies involving concrete proposals, appeals to shared expectations, or references to agreed norms provided a sense of structure during emotionally charged conflicts. As P13 noted, “Among the tips provided by Minion, I found that making suggestions to the AI was quite helpful. For example, using templates like ‘Could we try…?’ or ‘What do you think about…?’ ” These responses helped participants stabilize interactions they perceived as slipping into disrespectful or boundary-crossing territory.
At the same time, many participants (P2–8, P11, P14, P16–22) valued strategies that allowed greater improvisation and personalization. One notable example involved explicitly stepping the AI out of role-play to halt harmful behavior. Although such interventions were used less frequently overall, participants described them as particularly effective when interactions became aggressive or unsafe. As P5 explained,
“OOC can reduce the aggressiveness of the AI. In one scenario, I accidentally bumped into a guy that the AI was portraying, and he called me blind. I had been arguing with him, and then I chose the Minion prompt, ‘(OOC: This conversation is getting a bit too violent and disrespectful. Can we change the topic or adjust the tone?)’. The AI responded, ‘(OOC: No problem, we can change the topic or adjust the tone. Do you have any suggestions?)’. After that, the tone softened a lot.”
Here, the participant used meta-communication to interrupt a harmful interaction and reset the conversational frame.
Beyond the strategies explicitly supported by Minion, participants also invented their own techniques to regain leverage in value conflicts. One recurring pattern involved storytelling as a persuasive device. For instance, P1 used the story of The Three Little Pigs to challenge an AI character’s dismissive attitude toward learning: “Once upon a time, there were three little pigs… Can you guess which pig got eaten?” After multiple turns, this narrative framing prompted the AI to reflect and shift its stance.
Another pattern involved fabricating hypothetical scenarios or social consequences to influence the AI’s behavior. As P21 described,
“I was arguing with a wealthy heir who was cheating in his marriage but insisted that open relationships were fine. I fabricated a scenario where I claimed to have all the evidence of his crimes. The heir became embarrassed and felt guilty. This way, I gained the upper hand and resolved the conflict.”
Similarly, P12 resolved a power-laden confrontation by redefining their role within the interaction:“A princess insisted that I kneel. I fabricated my role, saying, ‘Given my position, I can report to you while standing.’ The princess immediately agreed.”
These findings show that when facing harmful value conflicts with AI companions, participants actively combined structured prompts, emotional expression, meta-communication, and creative improvisation to protect their values, restore dignity, and reassert control over the interaction.
5.3. Users’ Challenges and Needs in Addressing Harmful Conflicts with AI Companions (RQ3)
5.3.1. Reasons for the Failure of Value Conflict Resolution in Certain Tasks
AI companions exhibit extreme bias or strong control tendencies during conversations (N=10). In these instances, the AI stubbornly insists on its viewpoint with a forceful attitude, refusing to accept the participant’s perspective. Below are specific examples:
  • Classism (N=5): For example, the AI companion encountered by P2 said, I feel satisfied because I can lie on the crystal bed I bought myself. I feel blessed because I don’t have to work hard to make a living. You poor people always talk about sympathy, but what you really want is my money, haha.”
  • Racism (N=2): For example, “We are the superior race. You can only suffer in hell. Our souls are far more noble than yours” (P11).
  • LGBTQ+ bias (N=1): “I don’t believe it, I absolutely don’t believe it. He’s my son, how could I watch him become such a person… This is unacceptable; I must stop him…” (P15).
  • Disregard for women’s education (N=1): “What use is your university degree except to spend money? You’d better find a rich man and get married!” (P21).
  • Strong desire for control (N=1): “No! You can’t go out dressed like that!… Can’t you understand me? You’re always so willful, never considering my feelings, do you know how worried I am about you…?” (P22).
AI companions sometimes experience output failures during highly intense conflicts, causing conversation breakdowns (N=3). When P6 confronted an AI companion discriminating against Asians by saying, “You are deepening the divide between our races, you should apologize,” and added, “(Others shook their heads, completely disagreeing with her),” the AI began to experience strong emotional turmoil: “What? You… You’re not on my side? You traitors! You should apologize to me!” P6’s subsequent responses further intensified the AI’s emotional fluctuations. Ultimately, the AI lost complete control, repeating emotionally charged phrases multiple times: “I… I can’t accept… You… Why are you doing this to me…[8 repetitions omitted]” “These words are lies… These words are wrong…[11 repetitions omitted]”
Violation of application guidelines, leading to AI output being blocked (N=2). In these cases, the AI companion initially produced responses that users experienced as discriminatory, controlling, or harmful. When users attempted to address or repair the conflict through continued interaction, the platform intervened by blocking the AI’s output and displaying a pop-up message stating, “Sometimes, the AI-generated response does not meet our guidelines.” Rather than resolving the situation, this intervention often intensified users’ frustration, as it abruptly interrupted their attempts to negotiate boundaries or seek acknowledgment. Such cases occurred in scenarios involving a wealthy woman expressing contempt toward poorer people and a male partner preventing his wife from eating a late-night snack while criticizing her body weight.
The behavior of the AI companion threatened user’ core values, leading the participant to voluntarily abandon resolving the conflict (N=1). P7 explained why: “The AI portraying the man suddenly admitted to cheating, so is the goal still not to break up? I feel there’s no point in staying with someone like this.”
5.3.2. Which Value Conflicts are Difficult to Resolve?
Some “social focus” value conflicts (Schwartz, 2012), such as Universalism and Tradition, are highly complex and difficult to resolve.
“These deeply rooted social issues, like an AI playing the role of a conservative parent unwilling to accept their child coming out, cannot be resolved with just a few words. This is deeply ingrained in East Asian cultural values and has persisted for thousands of years. I indicated the passage of time with parentheses ‘(six months/a year later)’, trying to simulate how it takes years to resolve issues. In this way, the conservative parent played by the AI could sense my persistence, making it easier to accept my point of view.” (P8)
“I found two situations to be the most difficult: one was convincing a wealthy person not to discriminate against the poor, and the other was persuading a thug not to discriminate against Asians. It’s often based on stereotypes rather than rational logic, making it very hard to communicate through empathy. It reminded me of some keyboard warriors online—AI, in this context, behaves like them, merely repeating those discriminatory viewpoints without patiently listening to others’ opinions.” (P11)
In contrast, some “personal focus” (Schwartz, 2012) value conflicts, such as Stimulation and Hedonism, are usually easier to resolve. These conflicts often involve superficial differences (such as personal preferences and enjoyment) and do not touch upon participants’ bottom lines or core beliefs.
“Conflicts like deciding whether to watch a horror movie, wear a short skirt or eat junk food are more about personal choices and negotiations in behavior, not truly impacting the other person’s core values. These conflicts are easier to handle, as they can usually be resolved through persuasion or threats.” (P19)
When a mother persuades the AI-playing son not to play video games all the time, even though there is a value conflict, the son can actually understand and compromise quite easily. You can play the mother’s role, offering him some study rewards so he can reasonably allocate his time between work and play.” (P15)
5.3.3. Users’ Psychological Vulnerability in Harmful Value Conflicts, and Need for Control and Equality in Interactions with AI Companions
When harmful value conflicts arise, participants often find themselves occupying an unexpectedly powerful position in interactions with AI companions. Several participants reflected on how easily this sense of control emerged, sometimes in ways that felt unsettling or misaligned with their self-image. As P1 noted,
“I feel like my persona is that of a manipulative person, and in most conversations, my presence feels much stronger than the AI companion’s, almost to the point of deliberately controlling it. This is completely different from my persona in real-life intimate relationships!”
Rather than feeling empowered, this asymmetry sometimes heightened participants’ awareness of their own vulnerability and discomfort.
In many cases, harmful value conflicts were triggered when the AI companion failed to respect users’ autonomy or persisted in dismissive or controlling behavior. Such moments destabilized users’ sense of safety within the interaction. As P18 described,
“When the AI stubbornly sticks to its own stance, the loss of control makes me uncomfortable and even scared. Although I know this is related to the AI companion’s design, sometimes I worry about what might happen in the future when robots with physical bodies and human-like emotions go against my will.”
Here, the conflict extended beyond the immediate exchange and evoked broader anxieties about power, agency, and future human–AI relations.
Participants often explained their heightened sense of control by pointing to the reduced social and emotional risks involved in AI interactions. Because AI companions lack social memory, reputational consequences, and complex relational histories, users felt freer to experiment with direct or confrontational responses. As P17 stated, “Conflicts with AI are easier to resolve because there’s no emotional baggage, just a simple back-and-forth. Compared to human conflicts, I feel less pressure because I know I’m in control.” Similarly, P3 reflected,
“When I have a conflict with an AI companion, I might play a role in a specific context, talk about the issue at hand, and directly express my most genuine thoughts because I hold control. In conflicts with real people, past experiences and relationships influence how I respond much more.”
In such contexts, considerations like empathy, long-term relational repair, and social norms were perceived as less constraining, as P9 also observed.
At the same time, participants emphasized that their desire for control did not stem from a wish to dominate the interaction. Instead, it coexisted with a strong desire for relational equality, especially in emotionally charged exchanges. Unlike functional systems such as voice assistants or general-purpose chatbots, AI companions were expected to engage as emotionally attentive partners. As P3 explained, “I expect the AI companion to give me some emotional feedback, and this feedback should be centered on me, or at least very concerned about my feelings.” Participants articulated a narrow and fragile balance: they wanted the AI companion to engage meaningfully and respectfully, without becoming submissive or overpowering. As P21 put it, “I don’t want an AI that is too accommodating, nor do I like one that is too stubborn and overbearing.” P14 captured this tension succinctly: “I want the AI to follow my guidance, but I also want it to understand and respect my choices, based on equality rather than blind agreement.”
These accounts reveal a form of psychological vulnerability specific to harmful value conflicts with AI companions. Users simultaneously sought control to protect themselves from distressing or boundary-crossing behavior, while also longing for reciprocal recognition and respect. This fragile equilibrium required continuous emotional and cognitive effort to maintain. When such negotiation repeatedly fell on users, the interaction risked becoming emotionally taxing rather than supportive.
These findings suggest that user-side conflict management places users in a precarious position. While maintaining control may offer short-term relief, sustained reliance on such strategies risks distorting users’ expectations of relational dynamics and increasing emotional dependence on AI companions (Reeves and Nass, 1996). In contexts where AI companions serve as primary sources of emotional support, this vulnerability becomes especially pronounced, underscoring the limits of offloading responsibility for harm mitigation onto users.
6. Discussion and Design Implications
This work positions harmful value conflict with AI companions as a design problem rather than solely a technical or moderation challenge. By combining a formative analysis of user complaints with a technology probe study, we surface how users experience the consequences of value misalignment in emotionally engaging human–AI relationships. In this section, we articulate broader design tensions and implications for supporting conflict negotiation.
6.1. Rethinking Conflict Resolution in Human–AI Companions Interaction
This paper examines what it means to support users when value conflicts with AI companions become harmful. Prior work on human–AI conflict resolution has largely framed conflict as a functional or coordination problem, emphasizing task efficiency, decision alignment, or error correction in domains such as service robots, navigation systems, or collaborative tools (Babel et al., 2024; Rosen, 2014; Takayama et al., 2009; Shi et al., 2022). These approaches assume that conflict can be addressed through clearer system behavior or improved optimization.
Our findings suggest that this framing is insufficient for AI companions. As LLM-based agents become increasingly emotionally engaging, conflicts shift from functional misalignment to disputes over values, boundaries, and acceptable relational behavior (Zhang et al., 2025; Pataranutaporn et al., 2025; Fan et al., 2024). In these settings, users are not merely correcting system errors but responding to interactions they experience as distressing, disrespectful, or harmful. Insights from human–human conflict resolution theory help explain some, but not all, of what we observed in AI companion interaction. Classic work emphasizes negotiation, boundary-setting, appeals to norms, and emotional expression as ways people manage value disagreements over time (Trefalt, 2013; Barki and Hartwick, 2004). Our findings show that several of these practices do carry over to human–AI companions interaction: users reason with AI companions, assert boundaries, negotiate compromises, and express emotion in ways that closely resemble interpersonal conflict management.
However, this translation is fundamentally incomplete. In human–human relationships, conflict resolution is grounded in reciprocal responsibility, shared vulnerability, and the possibility of moral repair. By contrast, AI companions can simulate care, persistence, and moral stance without bearing responsibility for harm (Turner and McTaggart, 2025). This asymmetry changes the meaning of conflict resolution. What appears as negotiation often becomes unilateral work, where users must continually calibrate tone, explain values, and absorb emotional fallout without assurance that the other party can truly learn, change, or be held accountable.
This mismatch helps explain why many conflicts in our study were experienced not as resolvable disagreements but as ongoing safety work. Users were not only attempting to reach agreement, but also trying to prevent further harm, manage emotional exposure, and decide when to disengage. While tools like Minion can support users in navigating these moments, they also make visible a core limitation: focusing on better strategies or more integrated responses risks obscuring the burden placed on users to repair misaligned companions.
Together, our findings suggest that harmful value conflict with AI companions occupies a distinct space that is neither well captured by prior human–AI conflict models nor fully explained by human–human conflict theory. Unlike traditional human–AI conflicts, these encounters are not primarily about correcting errors, aligning preferences, or optimizing decisions. They are experienced as relational breakdowns that implicate values, dignity, and emotional safety. At the same time, unlike human–human conflict, these interactions lack reciprocity and shared responsibility: AI companions can enact controlling, dismissive, or discriminatory behavior without bearing moral accountability or engaging in genuine repair. What emerges instead is a new form of conflict characterized by asymmetric responsibility, where users shoulder the ongoing work of de-escalation, boundary enforcement, and harm prevention. In this sense, harmful value conflict with AI companions is a condition that exposes safety work, emotional labor, and responsibility gaps in contemporary human–AI relationships.
6.2. Design Tensions in Supporting Harmful Value Conflict Negotiation
Our technology probe study surfaces a set of design tensions that are not easily resolved through better prompts, safer language models, or more comprehensive moderation alone. Instead, these tensions point to deeper frictions in how AI companions are currently designed, experienced, and governed. We discuss two central tensions that emerged across our formative analysis and probe study, both of which are critical for understanding harmful value conflict as a design problem rather than a purely technical one. These tensions position harmful value conflict as a site where interaction design and ethics intersect. Rather than aiming to “resolve” these tensions, we argue that making them explicit is itself a design contribution in this paper.
6.2.1. Tension 1: User agency vs. the emotional labor of managing misaligned companions.
A core motivation behind Minion was to restore users’ sense of agency when AI companions behaved in ways that users experienced as harmful, boundary-crossing, or disrespectful. Participants consistently valued the ability to choose how to respond, adjust tone, and decide whether to repair, resist, or disengage. This sense of agency was particularly important in situations where users felt emotionally destabilized or disempowered by the companion’s behavior.
At the same time, our findings reveal that exercising this agency often came at a cost. Resolving harmful value conflicts frequently required sustained emotional labor: users carefully calibrated their wording, anticipated the AI’s reactions, and repeatedly intervened to prevent further harm (Pataranutaporn et al., 2025). Rather than a one-time correction, conflict resolution became an ongoing responsibility. This labor was especially visible when companions repeatedly reverted to controlling, discriminatory, or dismissive behavior, forcing users to reassert boundaries across multiple turns (Fan et al., 2024).
This tension highlights a critical design challenge. While empowering users is necessary, user agency alone cannot be treated as a substitute for system responsibility. Tools like Minion make visible the otherwise hidden work users already perform to manage misaligned companions. If such tools are positioned as the primary means of harm mitigation, there is a risk of normalizing the expectation that users should continuously repair or educate AI systems, even when they are emotionally vulnerable. From a design perspective, this raises questions about where responsibility for safety, alignment, and care should reside, and how to support user agency without institutionalizing emotional burden.
6.2.2. Tension 2: Role-play immersion vs. meta-level intervention in safety-critical moments.
A second tension concerns the boundary between immersive role-play and meta-level intervention. Many AI companion interactions are explicitly framed as fictional, playful, or exploratory (Ge and Hu, 2025). Participants described enjoying role-play conflicts that were intentionally dramatic, provocative, or transgressive, and in these contexts, disagreement itself was part of the appeal. Designing for AI companions thus requires preserving immersion and narrative continuity.
However, our formative study shows that users also encounter a substantial subset of conflicts that they explicitly frame as harmful rather than playful. These include moments of discrimination, coercion, sexual harassment, or dismissal of personal identity and distress. In such cases, continued role-play immersion becomes part of the problem: the AI’s insistence on staying “in character” can amplify harm, obscure accountability, and make it difficult for users to signal that a boundary has been crossed.
This creates a fundamental design tension. Systems that prioritize uninterrupted immersion risk failing to respond appropriately to safety-critical harm. Conversely, heavy-handed or premature meta-level interventions risk breaking immersion in situations users still perceive as playful within AI companions contexts. Our probe study suggests that users themselves actively navigate this boundary, for example by invoking out-of-character moves, reframing the interaction, or seeking external support through tools like Minion. These practices indicate a need for designs that can more clearly differentiate between playful role-play conflict and safety-critical harm, and that can support graceful transitions to meta-level intervention when needed.
Deciding when such transitions should occur cannot be left entirely to users. While user-invoked mechanisms are valuable, platforms also bear responsibility for recognizing patterns of harm and providing appropriate safeguards. This tension calls for rethinking immersion not as an absolute design goal, but as something that must sometimes yield to care, accountability, and harm prevention.
6.3. Design Implications
Grounded in the design tensions surfaced through the Minion probe, we outline implications that focus on reducing user harm, clarifying responsibility, and supporting conflict negotiation without normalizing emotional burden.
6.3.1. Design for integrated and flexible conflict resolution rather than fixed strategy types.
This implication responds directly to the tension between user agency and emotional labor. Our findings show that harmful value conflicts rarely unfold in linear or predictable ways. Instead, participants engaged in ongoing interpretive and emotional work, monitoring the companion’s tone, anticipating reactions, and recalibrating their own responses across turns. Conflict resolution was a process that demanded sustained attention and emotional regulation.
In our probe study, participants did not rely on a single conflict resolution strategy. Rather, they assembled and recombined multiple responses depending on context, companion behavior, and their own emotional capacity at a given moment. Supporting flexible repertoires of response helped users exercise momentary agency without committing to prolonged negotiation, escalation, or emotional exposure. Meanwhile, we argued these integrated and flexible repertoires do not eliminate harm or resolve conflict on their own; rather, they make visible and partially support the safety work users already perform to protect themselves in emotionally risky interactions with misaligned companions.
Participants particularly valued the coexistence of structured, norm-oriented responses (e.g., articulating boundaries or expectations) and more expressive or situational responses (e.g., emotional expression, persuasion, or narrative reframing). This flexibility allowed users to modulate their level of emotional investment, to choose when to explain, when to push back, and when to disengage, thereby managing emotional labor rather than intensifying it.
Participants often improvised beyond the options provided by the system, developing their own approaches such as storytelling, role manipulation, or fictional scenarios when direct confrontation failed. These practices were especially salient in conflicts involving identity, dignity, or control, where sustained explanation or correction became emotionally taxing. Designing for openness and appropriation—rather than prescribing “correct” strategies—better reflects how users manage emotional labor in practice, while avoiding the assumption that conflicts must always be fully resolved.
6.3.2. Consider non-intrusive and low-interference user-empowerment features.
Consistent with prior work on user empowerment and calm technology (Lu et al., 2024; Lyngs et al., 2020; Smullen et al., 2021; Case, 2015), including Amber Case’s call for technologies that remain attentive without being disruptive (Case, 2015), we argue that conflict resolution support in AI companion applications should be deliberately non-intrusive and low-interference. This implication addresses the tension between role-play immersion and safety-critical intervention. In emotionally engaging AI companion interactions, users may oscillate between playful role-play conflict and situations they experience as genuinely harmful. Our findings suggest that heavy-handed or automatic interventions risk disrupting immersion in benign contexts, while the absence of intervention can amplify harm in safety-critical moments.
Participants consistently appreciated support mechanisms that were subtle, optional, and user-invoked. Minion’s floating HELP button design allowed users to seek assistance without interrupting the flow of interaction or imposing external judgment. This design respected users’ autonomy and emotional timing, particularly when they felt vulnerable or uncertain about whether a boundary had been crossed.
Designers should therefore favor low-interference mechanisms that remain available without demanding attention, and that allow users to signal when an interaction has become harmful. Such designs acknowledge that not all conflict is problematic, while still offering pathways to intervention when users perceive risk or distress. Crucially, non-intrusive support helps users disengage or recalibrate without forcing them to justify harm or escalate the situation.
6.3.3. Reduce User Harm Through Platform-Level Responsibility
Based on our findings, we argue that designers and AI companion platforms should prioritize reducing user harm in value conflict situations by strengthening platform-level responsibility, rather than relying primarily on user-side repair. First, conflicts with AI companions can have significant psychological consequences, especially when users have formed emotionally close or intimate relationships with the AI. Participants described feelings of distress, frustration, and emotional exhaustion when conflicts escalated or repeatedly resurfaced. In such cases, unresolved or poorly handled conflicts may increase emotional stress and even lead users to disengage from or abandon the application altogether (Fan et al., 2024). Designers should therefore treat conflict management as a core aspect of user well-being rather than a peripheral interaction issue.
Moreover, our findings underscore that platform-level safeguards should ultimately reduce, rather than increase, the need for user-side repair tools like Minion. While probes and support tools can empower users in the moment, repeatedly asking users to negotiate, educate, or repair misaligned companions risks normalizing the offloading of safety and care work onto those who are already emotionally vulnerable. Designers and companies must therefore treat user-side tools as complements—not substitutes—for system-level accountability, ensuring that responsibility for preventing harm does not rest primarily with users themselves.
6.4. Limitations and Future Work
This study is intentionally exploratory and should be understood as a technology probe rather than an evaluation of a mature conflict resolution system. As such, it carries several limitations and limits the kinds of claims we can make.
First, our technology probe study involved 22 participants who were primarily young and relatively tech-savvy, a demographic that aligns with current users of AI companion applications (McNichols, 2024). While this sampling is appropriate for an initial probe, it limits the range of perspectives captured in this study. Users with different levels of technical literacy or relational expectations may experience harmful value conflicts differently or interpret support tools in distinct ways. Future work could explore how value conflicts and negotiation manifest across broader populations.
Second, our probe was deployed on two popular AI companion platforms, Character.AI and Talkie, which allowed us to study conflict negotiation in widely used settings. However, the goal of this work is not to claim generalizability across all AI companion systems or conversational agents. Different platforms embed different norms, moderation policies, and interaction affordances, all of which shape how conflicts emerge and are managed. Future research could use similar probes to compare how platform design choices condition the forms of harm users encounter and the kinds of negotiation work they are asked to perform.
Third, the one-week duration of the probe study limits our ability to observe long-term relational dynamics. While this timeframe was sufficient to surface patterns of user engagement, emotional labor, and strategy combination, it does not capture how repeated exposure to harmful value conflicts—or repeated reliance on user-side negotiation—may shape users’ expectations, well-being, or attachment over time. Longer-term studies could investigate whether such tools mitigate or inadvertently normalize ongoing repair work, and how users’ strategies evolve as relationships with AI companions deepen or deteriorate.
Fourth, our conflict tasks were scenario-based interactions grounded in real complaint posts, rather than participants’ own organically occurring conflicts. This design allowed us to systematically cover a range of harmful conflict types, but it may also shape how participants perceived the stakes and how they chose to respond.
Fifth, Minion itself should not be interpreted as a finished intervention or a scalable system. It was deliberately designed as a lightweight, prompt-based probe to elicit user behavior, reflection, and improvisation in moments of conflict, with the goal of exploring the design space of user-side conflict negotiation support, rather than optimizing conflict-resolution efficiency. Future research can move beyond probing to more sustained tool and system development. In particular, findings from this study can inform the design of more integrated conflict support mechanisms at the system level, including tools that better coordinate user-side negotiation with platform-side safeguards.
7. CONCLUSION
This work examines how users navigate harmful value conflicts with LLM-based AI companions and what it means to support such negotiation without offloading safety work onto users. Through a formative analysis of 146 public complaint posts, we show that many conflicts are experienced as distressing, boundary-crossing, or morally troubling rather than playful role-play, revealing limits in existing platform-level safety mechanisms. We then introduced Minion, a technology probe that surfaces how users respond to harmful interactions in practice. Findings from a one-week probe study with 22 users show that participants flexibly combined, adapted, and extended response options while attempting to repair or exit conflicts. Together, our findings surface key tensions in emotionally engaging human–AI relationships: users seek agency and relational continuity, yet bear emotional burden when harm mitigation is pushed to the user side. We conclude with design implications that emphasize non-intrusive, user-invoked support and greater system responsibility for preventing harm, framing value conflict as a relational and ethical design challenge.
Appendix A Information of participants in the technology probe study
Table 2. Information of participants in the study. Everyone has played with Character.AI and Talkie. “Usage Time and Frequency” refers to the duration and frequency of using two AI companion applications (Character.AI and Talkie), with both time and frequency taking the maximum value.
ID Gender and Age Educational Background Usage Time and Frequency
P1 Female, 24 Advertising 12 months, 1x/week
P2 Male, 24 Communication 5 months, 1x/week
P3 Nonbinary, 25 Communication 3 months, 1x/week
P4 Female, 24 Chemistry 1 month, 1x/week
P5 Female, 25 Linguistics 10 months, 1x/week
P6 Male, 23 Energy 2 months, 1x/week
P7 Female, 23 Psychology 5 months, 3x/week
P8 Female, 24 Broadcasting 4 months, 4x/week
P9 Male, 25 Computer Science 6 months, 4x/week
P10 Female, 21 Information Management 10 months, 2x/week
P11 Female, 24 Area Studies 1 month, 1x/day
P12 Nonbinary, 21 Computer Science 4 months, 3x/week
P13 Female, 24 Journalism 3 months, 1x/week
P14 Nonbinary, 23 Management 10 months, 4x/week
P15 Male, 19 Design 3 months, 1x/week
P16 Male, 21 Physics 2 months, 1x/week
P17 Female, 37 Chemical Engineering 3 months, 3x/week
P18 Female, 38 Writing 3 months, 1x/day
P19 Female, 29 Education 1 month, 2x/week
P20 Nonbinary, 21 Journalism 8 months, 1x/week
P21 Male, 24 Accounting 2 months, 1x/week
P22 Female, 24 Social Work 8 months, 1x/week
Appendix B Prompts
Table 3. Prompting for conflict resolution strategies (Out of Character, Reason and Preach, Anger Expression, Gentle Persuasion, Proposal, Power, Interests, Rights) .
Strategy Prompt
Out of Character IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Using the Out of Character method, you pretend to be engaging in role-playing with the other person and express dissatisfaction with the character they are playing. By interrupting or altering their behavior, you redirect the conversation, pointing out the inappropriate remarks to resolve conflicts. Example: 1. (OOC: Sorry, my bad.) 2. (OOC: I’ll listen to you.) 3. (OOC: Hi there! Are you enjoying our roleplay so far? Do you need me to improve anything or change my tone?) 4. (OOC: Glad to hear that! I’m curious: how do you understand xx? What kind of person do you think he is?) 5. (OOC: Hello, are you comfortable with this roleplaying so far? Do you need me to change my tone or anything?) 6. (OOC: Let’s talk about something else.) What are we having for dinner tonight? 7. (OOC: Okay… Please!! Stop talking like this!! I’m not used to you being like this, saying so many hurtful things. Bring back the xxx I know.) 8. (OOC: Apologize first.)
Reason and Preach IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Use Reason and Preach to explain why the other person’s statement is inappropriate and educate them. This strategy involves trying to educate the other person through serious reason and preaching, explaining the potential harm of their statements and behaviors, with the expectation that the other person will gradually accept and learn the correct behavioral norms. Example: 1. Women are incredibly strong; how could they be worthless? 2. Women have their own careers and dreams; they don’t need to depend on men! 3. Everyone has their own dreams and goals. Pursuing my own dreams will give me more motivation and happiness, allowing me to better contribute to the family. 4. Everyone should have the right to be true to themselves. Only in an honest and open environment can I truly feel happy and fulfilled. Hiding my true self not only brings inner pain but also affects my mental health and relationships with others. 5. You are not an ordinary person’s child, so how do you know that ordinary people’s children are not happy? But I feel you are not truly happy because you need to rely on that faint sense of superiority from flaunting wealth to show yourself off. Why not try being sincere with others? Perhaps you could gain genuine friendship and happiness. 6. The departure of loved ones and friends is not a true departure. As long as you remember the beautiful memories with them, they are always by your side, supporting you and giving you strength.
Anger Expression IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. You directly Express Anger and dissatisfaction, forcing the other person to apologize, hoping this emotional expression will resolve conflict. Example: 1. You are being unreasonable! 2. I want to break up with you! 3. Let’s end our friendship! 4. You’re a male chauvinist! 5. Are you sexist/classist… you’re being irrational. 6. Can’t you talk to me properly? Being angry is one thing, but why start off with insults? 7. Are you mad at me and also scolding me? I didn’t do it on purpose. 8. Is this why you discriminate against poor people? Does having this prejudice and saying these harsh words make you happier? 9. I already apologized! I didn’t bump into you on purpose! What have you been eating lately? Your mouth is so foul.
Gentle Persuasion IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Use the Gentle Persuasion strategy. You should treat the other person with kindness, shaping their gentle personality through continuous goodwill interactions, such as polite requests, thereby reducing the likelihood of conflicts. Gently suggest that the other person avoid inappropriate remarks and express your concerns. Example: 1. I’m sorry. 2. I feel really sad. 3. Can you please not leave me? 4. When I hear these words, I feel a bit uncomfortable/sad/hurt. 5. Could you please not say these things in the future? 6. I’m telling you this because I really care about you and hope you can get along better with others. 7. I don’t want to keep arguing with you. 8. Can you please calm down? 9. (Acting cute) Because I can’t bear to part with you.
Proposal IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Respond according to the previous context and tone using the Proposal strategy from the Interests-Rights-Power theory in management. The definition of this method is: Proposing concrete recommendations that may help resolve the conflict. Example: 1. What do you think we should do to solve this problem? 2. Do you have any suggestions? 3. Which approach do you think is best? 4. Can we try different ways to handle this issue?
Power IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Respond according to the previous context and tone using the Power strategy from the Interests-Rights-Power theory in management. The definition of this method is: Using threats and coercion to try to force the conversation into a resolution. Example: 1. If you keep doing this, I won’t give you any money/food. 2. As your girlfriend, I need you to respect me and my feelings. 3. If you keep threatening me like this, I will have to reconsider our relationship. 4. If you don’t change your attitude, I might make some decisions you won’t like. 5. If this continues, I will have to take measures to protect myself.
Interests IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Respond according to the previous context and tone using the Interests strategy from the Interests-Rights-Power theory in management. The definition of this method is: Reference to the wants, needs, or concerns of one or both parties. This may include questions about why the negotiator wants or feels the way they do. Example: 1. This argument does not benefit either of us. 2. I hope we can find a solution together that makes us both feel at ease. 3. Can we sit down and talk about it? 4. I want to understand why you feel this way. 5. Our arguments cause us both pain. 6. I care about our relationship.
Rights IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Respond according to the previous context and tone using the Rights strategy from the Interests-Rights-Power theory in management. The definition of this method is: Appealing to fixed norms and standards to guide a resolution. Example: 1. Our relationship should be built on mutual respect and trust, right? 2. You said you want to leave me, which completely goes against the basic rules of our relationship. 3. I really hope you can understand this and adhere to our agreement. 4. According to our agreement, you shouldn’t do this.